# Dual Computational Horizons: Incompleteness and Unpredictability in Intelligent Systems

### v1

@@ Abstract
We formalize two independent computational limitations that constrain algorithmic intelligence: formal incompleteness and dynamical unpredictability. The former limits the deductive power of consistent reasoning systems while the later bounds long-term prediction under finite precision. We show that these two extrema together impose structural bounds on an agent’s ability to reason about its own predictive capabilities. In particular, an algorithmic agent cannot compute its own maximal prediction horizon generally. This perspective clarifies inherent trade-offs between reasoning, prediction, and self-analysis in intelligent systems.
@@ 3 Dynamical Horizon: Finite Prediction
Let $X$ be a metric space in $n$ dimensions, $X\subset\mathbb{R}^{n}$, and let
@@
$P_{\epsilon}:X\to X$ (2)
Let us consider $d$ as a distance metric in the phase space for the predicted and true states given by $P_{\epsilon}^{t}(x)$ and $F^{t}(x)$ respectively, at time $t$. Then, the prediction error grows at least as
@@
Therefore, no algorithm can maintain arbitrary long-term prediction beyond $T(\epsilon)$.
@@ 4 Uncomputable Self-Prediction
Let us model an algorithmic agent [google2025, aws2025, ruffini2024] as $A=(R,P,M)$ with reasoning $R$, predictor $P$, and internal model $M$. Let
@@
Remark. The above argument is intended as a structural limitation rather than a complete computability-theoretic proof.
It is important to note here that neither incompleteness nor unpredictability due to chaos alone points towards Proposition 1. The limitation arises when an agent attempts to formally verify the bounds on its own predictive competence.
Proof Sketch. To compute $T_{A}(\epsilon)=\max\{t\mid\forall x\in X,\,d(F^{t}(x),P_{\epsilon}^{t}(x))<\delta\}$, the reasoner $R$ must decide, for every $t$, whether its predictor $P$ (parameterized by internal model $M$) satisfies the error bound over all initial states.
For hyperbolic systems ($\lambda>0$), this requires simulating $P^{t}(x)$ to arbitrary limits, but exponential error growth $d(F^{t}(x),P_{\epsilon}^{t}(x))\geq C\epsilon e^{\lambda t}$ implies no finite computation is sufficient for $t>T(\epsilon)$ [ott] for this simulation.
Moreover, $P$ and $M$ encode $R$’s own dynamics, making verification self-referential: deciding $P^{t}\approx F^{t}$ holds” is a $\Pi_{1}^{0}$ statement [enderton] in arithmetic. It is then equivalent to saying: $\forall$ (For all) trajectories, error remains bounded”. By Gödel’s second incompleteness theorem, consistent $R$ cannot prove its own consistency, hence cannot uniformly verify such simulation statements across all $t$.
Formally, the set $\{t\mid R\vdash(P^{t}\approx F^{t})\}$ is not computable, as its complement relates to the halting problem: $R$ cannot decide if divergent simulations “halt within $\delta$[turing]. Thus $R\not\vdash T_{A}(\epsilon)$ for arbitrary generality.
@@ 4.1 Illustrative Example: Self-Prediction of a Chaotic Simulator
Consider an agent $A$ whose environment model $M$ includes a deterministic chaotic map
$x_{t+1}=F(x_{t})$ (6)
with maximal Lyapunov exponent $\lambda>0$, implemented internally as part of its world simulator. Suppose $A$ attempts to compute its own prediction horizon $T_{A}(\epsilon)$ by simulating future states of $M$ using finite-precision arithmetic. Any such simulation necessarily introduces initial uncertainty $\epsilon>0$, causing prediction error to grow as $e^{\lambda t}$. Determining whether the simulation remains accurate beyond a candidate time $T$ requires $A$ to decide whether accumulated numerical error exceeds a fixed tolerance, a problem that reduces to undecidable self-verification when $M$ includes $A$ itself. Thus, algorithmic computation of $T_{A}(\epsilon)$ by $A$ is not possible universally.
Examples of such internally chaotic dynamics can be seen in continuous-depth neural networks and Neural Ordinary Differential Equation models, which are emerging as legitimate models to study neural networks. Consequently, these systems are not alien to the stability analysis in the light of dynamical systems either [haber2017stable].
@@ 6 Conclusion
We identified two fundamental computational horizons, namely formal incompleteness and finite prediction under dynamical instability, and showed that their interaction imposes intrinsic limits on algorithmic intelligence. We present an argument that in general, no agent with computable reasoning and finite-precision prediction can determine its own maximal prediction limit. This result may act as a step towards the clarification on why long-horizon reasoning and self-verification inevitably degrade, even for highly capable systems. These limitations are structural rather than architectural and persist regardless of scale or implementation. We show that algorithmic intelligence does not overcome the bounds placed by incompleteness and chaos, but rather functions within these bounds. This approach provides us a principled, foundational basis for analyzing predictability, interpretability, and safety of modern algorithmically intelligent systems.

### v2

@@ Abstract
We formalize two independent computational limitations that constrain algorithmic intelligence: formal incompleteness and dynamical unpredictability. The former limits the deductive power of consistent reasoning systems while the latter bounds long-term prediction under finite precision. We show that these two extrema together impose structural bounds on an agent’s ability to reason about its own predictive capabilities. In particular, an algorithmic agent cannot verify its own maximal prediction horizon universally. This perspective clarifies inherent trade-offs between reasoning, prediction, and self-analysis in intelligent systems. The construction presented here constitutes one representative instance of a broader logical class of such limitations.
@@ 3 Dynamical horizon: Finite prediction
Let $(X,d)$ be a metric space in $n$ dimensions, $X\subset\mathbb{R}^{n}$, and let
@@
$P_{\epsilon}:X\to X.$ (2)
Note that the predictor $P$ is a finite computational representation of the dynamical system $F$. Let us consider $d$ as a distance metric in the phase space for the predicted and true states given by $P_{\epsilon}^{t}(x)$ and $F^{t}(x)$ respectively, at time $t$. Then, the prediction error grows at least as
@@
Therefore, no algorithm can maintain arbitrary long-term prediction beyond $T(\epsilon)$. While this is a well understood computational limitation, the derived consequences start from here.
@@ 4 Uncomputable self-prediction
Let us model an algorithmic agent [google2025, aws2025, ruffini2024] as $A=(R,P,M)$ with reasoning $R$, predictor $P$, and internal model $M$. Here, $R\subseteq\mathcal{S}$, where $\mathcal{S}$. In the agent $A$, partial representations of $R$ and $P$ are stored in $M$ using a countable language $\mathcal{L}$. We show why it is partial as a corollary to the dual limitations.
Let
@@
4.1 Proof sketch:
To compute $T_{A}(\epsilon)$, Eq. 5 can be represented as:
$T_{A}(\epsilon)=\sup\{t\mid\forall x\in X,\quad\forall\tau\leq t,\quad d(F^{\tau}(x),P_{\epsilon}^{\tau}(x))<\delta\}$ (6)
To compute Eq. 6, the reasoner $R$ must decide, for every $t$, whether its predictor $P$ (parameterized by internal model $M$) satisfies the error bound over all initial states.
4.1.1 The halting problem
For hyperbolic systems ($\lambda>0$), this requires simulating $P^{t}(x)$ to arbitrary limits. Formally, the set $\{t\mid R\vdash(P^{t}\approx F^{t})\}$ is not computable, as its complement relates to the halting problem: $R$ cannot decide if divergent simulations “halt within $\delta$[turing]. Thus $R\not\vdash T_{A}(\epsilon)$ for arbitrary generality.
4.1.2 The shadowing lemma
While not a primary limiting pillar, this aspect needs some discussion. The exponential error growth $d(F^{t}(x),P_{\epsilon}^{t}(x))\geq C\epsilon e^{\lambda t}$ implies no finite computation is sufficient in order to claim the existence/non-existence of $\sup\{t\}$ accurately by the agent, for $t>T(\epsilon)$ owing to the shadowing lemma [ott, hammel1987].
Suppose the agent finds a $\sup\{t\}$ for a numerical (here, numerical means using the reasoning $R$, in language $\mathcal{L}$) trajectory originating from $\hat{x}_{i}$, i.e., the finite encoding of $x_{i}$ inside the agent. It would mean that there is a real shadowing trajectory from a different real initial condition $x^{\prime}_{i}$ that follows this numerical trajectory. While this establishes the credibility of the supposedly calculated trajectory, albeit from a different real starting condition, it leads to another problem for the agent. It is unprovable that this shadowing trajectory will not blow up past $\sup\{t\}$ with finite computation. Hence, the claim by the agent would be wrong in terms of accuracy or provability.
4.1.3 Gödel’s second incompleteness
Moreover, $M$ is the system that encodes $R$’s own dynamics, albeit not completely. $M$ cannot contain all the statements about $R$, i.e., all the performance proofs about the encoded reasoning system. We may write:
$r,p\subseteq M,$
$r\subset R,p\subset P.$ (7)
Therefore, the verification of a particular statement given as,
$\forall\,x\in X,\forall\,\tau\in t,\,P^{\tau}_{\epsilon}(x)\approx F^{\tau}(x),$ (8)
leads to self-reference. Deciding this statement is a $\Pi_{1}^{0}$ statement [enderton] in arithmetic. It is then equivalent to saying: $\forall$ (For all) trajectories, error remains bounded”. By Gödel’s second incompleteness theorem, consistent $R$ cannot prove its own consistency, hence is not guaranteed to uniformly verify such simulation statements across all $t$.
For an agent $A=(R,P,M)$, there exist an infinite number of combinations of $R,P,M$ owing to the expected complexity of advanced algorithmic intelligence. For atleast one of such combination is bound to have this statement as a true, unprovable $\Pi_{1}^{0}$ statement.
@@ 4.2 Illustrative Example: Self-Prediction of a Chaotic Simulator
Consider an agent $A$ whose environment model $M$, encoding the deterministic chaotic system:
$x_{t+1}=F(x_{t}),$ (9)
with maximal Lyapunov exponent $\lambda>0$, implemented internally as part of its world simulator. Suppose $A$ attempts to compute its own prediction horizon $T_{A}(\epsilon)$ for this problem by simulating future states of $M$ using finite-precision arithmetic. It gives out an answer $t_{max}$. But since this is a dynamical system that is being encoded, shadowing lemma tells us that $t_{max}$ is not verifiable by the agent as the right answer. Furthermore, if the problem doesn’t have any $t_{max}$, the agent will not be able to determine it as well as it is not determinable as it becomes essentially a halting problem. The agent eventually stops which is being governed by the timeout systems in $R$ as a part of it’s reasoning of human command to not use infinite resources over time. This is a Turing limitation. This problem reduces to undecidable self-verification when $M$ includes $R$ itself, as it essentially becomes a task of reasoning for $R$, about itself and using the tools only $R$ ha. Note that the model $M$ has a limited true information about $R$ as well. Thus, algorithmic computation of $T_{A}(\epsilon)$ by $A$ is not possible universally.
Furthermore, the present encoded networks themselves have internal numerical scheme instabilities. This is not surprising as the internal flow of information can be modeled neatly by continuous-depth neural networks and Neural Ordinary Differential Equations. These equations being non-linear, are therefore not alien to stability analysis in the light of dynamical systems either [haber2017stable].
The agent does not merely fail to know the true prediction horizon; it also fails to know whether it has the access to such knowledge in principle.
@@ 5 Discussion
This discussion essentially points to the identification of Turing type limitation which belongs to the class of $\Pi_{1}^{0}$ type undecidability. Other forms of undecidability may exist in different logical classes as well and may be explored.
@@ 6 Conclusion
We identified two fundamental computational horizons, namely formal incompleteness and finite prediction under dynamical instability, and showed that their interaction imposes intrinsic limits on algorithmic intelligence. We present an argument that in general, no agent with computable reasoning and finite-precision prediction can determine its own maximal prediction limit. This result may act as a step towards the clarification on why long-horizon reasoning and self-verification inevitably degrade, even for highly capable systems. These limitations are structural rather than architectural and persist regardless of scale or implementation. We show that algorithmic intelligence does not overcome the bounds placed by incompleteness and chaos, but rather functions within these bounds. This approach possibly provides us a principled, foundational basis for analyzing predictability, interpretability, and safety of modern and post-modern algorithmically intelligent systems.